OpenStudy (anonymous):

How can I calculate the equation cos(xpi/6)=cos(xpi/4) for the smallest possible integer value?

OpenStudy (anonymous):

Like this: \[\cos(\frac{ x \pi }{ 6 })=\cos(\frac{ x \pi }{ 4 })\] Where x = smallest positive integer

OpenStudy (anonymous):

@asnaseer This is actually a follow up question to my previous one :P

OpenStudy (anonymous):

@wio Can you help me with this?

OpenStudy (anonymous):

@AravindG ?

OpenStudy (anonymous):

@ash2326 ?

OpenStudy (jdoe0001):

can you post a quick screenshot of the material?

OpenStudy (jdoe0001):

maybe there's something we're missing from the exercise

OpenStudy (anonymous):

No there isn't anything missing, is there something wrong with the equation? I know the answer, I just don't know how to get it.

OpenStudy (anonymous):

Oh there's a typo in the question - it should be smallest positive integer, not smallest possible integer

OpenStudy (asnaseer):

You could approach it as follows: \(\cos(\theta)=\cos(2\pi n+\theta)\) for \(n=1,2,...\) \(\therefore\cos(\pi x/6)=\cos(2\pi n+\pi x/6)=\cos(\pi(12n+x)/6))\) We therefore need to solve: \(\cos(\pi x/4)=\cos(\pi(12n+x)/6))\) i.e.: \(\pi x/4=\pi(12n+x)/6\) which leads to: \(x=24n\) So the smallest positive \(x=24\)